Calculate the uncertainty (in m) in the position of a 120 g baseball thrown at a velocity of 45m/s if the uncertainty in velocity is 2%.
Answer #1

Here:

velocity = 45 m/s

Δv = 2.0% of velocity

Δv = 2.0*45/100

Δv = 0.9 m/s

use Heisenberg uncertainty principle,

Δx*Δv = h/(4*pi*m)

Here:

Δv = 0.9 m/s

mass, m = 1.2*10^2 g

= 0.12 kg

Putting values,

Δx*(0.9) = (6.626*10^-34)/(4*3.14*0.12)

Δx*(0.9) = 4.396*10^-34

Δx = 4.885*10^-34 m

Answer: 4.9*10^-34 m